OpenStudy (anonymous):

You toss 5 coins. What is the probability that all 5 show heads? Can someone help please will fan and give medal

OpenStudy (welshfella):

the probability that on one toss you shoe heads = 1/2

OpenStudy (anonymous):



OpenStudy (anonymous):

form what fraction how many times could you get on all heads

OpenStudy (welshfella):

and for second and third etc its still 1/2 these are independent events so you multiply the probabilities

OpenStudy (anonymous):

5/5 ?

OpenStudy (arthur_ser):

\[\frac{ 1 }{ 2 }\times\frac{ 1 }{ 2 }\times\frac{ 1 }{ 2 }\times\frac{ 1 }{ 2 }\times\frac{ 1 }{ 2 }=(\frac{ 1 }{ 2 })^5\]

OpenStudy (welshfella):

arthur has it